# Global well-posedness for the generalized intermediate NLS with a nonvanishing condition at infinity

### v1

@@ 3. Local well-posedness
Let $\phi\in\mathcal{Z}^{2}_{\rho}$ satisfy $\mathcal{E}^{2}_{\rho}(\phi)\leq M_{0}^{2}$, and let $T_{*}=T_{*}(\delta,M_{0})>0$ be the constant given in Proposition 3.3. We divide the proof into several steps.
Step 1. We shall construct a local solution $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2}_{\rho})$ to (gINLS) with initial data $\phi$ satisfying
(3.23) $\big\|\mathcal{E}^{2}_{\rho}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim\mathcal{E}^{2}_{\rho}(\phi),$
and also show that if $\phi\in\mathcal{Z}^{k}_{\rho}$ for some integer $k\geq 3$, then $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{k}_{\rho})$ and it satisfies
(3.24) $\big\|\mathcal{E}^{k}_{\rho}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim_{k}\mathcal{E}^{k}_{\rho}(\phi).$
Let $\{u^{(\ell)}\}_{\ell\geq 0}$ be the sequence of solutions to the mollified problem (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ constructed in Proposition 3.1. We recall from Propositions 3.2 and 3.3 that $\{u^{(\ell)}\}_{\ell\geq 0}\subset C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$ and it is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$. Since $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$ is a complete metric space, there exists $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$ such that $u^{(\ell)}\to u$ strongly in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$. Moreover, (3.5) shows that the sequence $\{\partial_{x}u^{(\ell)}\}_{\ell\geq 0}$ is bounded in $L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. By the Banach-Alaoglu theorem, there exists a subsequence converging weakly-* to some $\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. Since $\partial_{x}u^{(\ell)}\to\partial_{x}u$ strongly in $C([0,T_{*}];L^{2}(\mathbb{R}))$, we have $\partial_{x}u=\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$, that is, $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{2}_{\rho})$. The bound (3.23) follows from (3.5) and weak-* lower semicontinuity of the $L^{\infty}_{T_{*}}H^{1}$ norm. The claim in the regular case $\phi\in\mathcal{Z}^{k}_{\rho}$ ($k\geq 3$) can be shown in the same way using (3.6).
It remains to show that the limit $u$ is a solution to (gINLS) with $u(0)=\phi$. The initial condition is clearly satisfied by the strong convergence in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$. To see that $u$ satisfies the integral equation
@@
$\lesssim 2^{-\ell}(1+\delta^{-1})\|\mathcal{E}^{2}_{\rho}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}^{\frac{3}{2}}+(1+\delta^{-1})\Big(1+\|\mathcal{E}^{1}_{\rho}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}+\|\mathcal{E}^{1}_{\rho}(u)\|_{L^{\infty}_{T_{*}}}\Big)\|d^{1}_{\rho}(P_{\leq\ell}u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{\delta,M_{0}}2^{-\ell}+\|d^{1}_{\rho}(u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
@@
where we have used Lemma 2.1, the estimate (2.11), the bounds (3.5), (3.23) and the strong convergence $u^{(\ell)}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$. In particular, we have the convergence of the Duhamel integrals
@@
This and the convergence of the linear part imply that $u$ satisfies (3.25).
@@
We conclude Step 1 by mentioning a bound of distance between two solutions of (gINLS), as follows. Let $\phi,\psi\in\mathcal{Z}^{2}_{\rho}$ with $\mathcal{E}^{2}_{\rho}(\phi),\mathcal{E}^{2}_{\rho}(\psi)\leq M_{0}^{2}$, and let $u,v\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2}_{\rho})$ be the corresponding solutions of (gINLS) constructed above. Then,
(3.26) $\big\|d^{1}_{\rho}(u,v)\big\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}_{\rho}(\phi,\psi).$
Indeed, this bound follows immediately by taking $\ell,m\to\infty$ in (3.15).
@@
Step 2. We will see that the following holds: Let $u,v\in C([0,T];\mathcal{Z}^{2}_{\rho})$ be two solutions to (gINLS) on some interval $[0,T]$. Assume that $\|\mathcal{E}^{2}_{\rho}(u)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ and $\|\mathcal{E}^{2}_{\rho}(v)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ for some $M_{1}>0$. Then, there exists $T_{1}=T_{1}(\delta,M_{1})>0$ such that for $0\leq t\leq\min\{T,T_{1}\}$ it holds that
$d^{1}_{\rho}(u(t),v(t))\lesssim(1+M_{1})d^{1}_{\rho}(u(0),v(0)).$
@@
In particular, the solution of the Cauchy problem associated with (gINLS) is unique in $C([0,T];\mathcal{Z}^{2}_{\rho})$.
@@
We remark that this step must be discussed independently of the previous Step 1 (in particular of the difference estimate (3.26)), since for uniqueness we have to consider general solutions which may differ from those constructed in Step 1. However, the claim can be proven along the same lines as Proposition 3.3 using the original equation (gINLS), (3.25) instead of the mollified one (3.1), (3.2). We omit the details.
@@
Step 3. We shall prove the continuity of the time evolution $t\mapsto u(t)$ in $\mathcal{Z}^{2}_{\rho}$ and that of the flow map $u_{0}\mapsto u$ in $C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$.
@@
We can prove the continuity of the evolution in $\mathcal{Z}^{k}_{\rho}$ for $\phi\in\mathcal{Z}^{k}_{\rho}$ in a similar manner; we omit the proof. In particular, this concludes the proof of Theorem 1.1. To prove the claim, we may adapt an argument due to Bona–Smith [Bona-Smith] or the idea of frequency envelope due to Tao [Tao-2001, Ifrim-Tataru, Burq] to the non-vanishing solutions in $\mathcal{Z}^{2}_{\rho}$.444 The abstract theorem in [Burq] does not seem to apply directly to our problem, since $\mathcal{Z}^{2}_{\rho}$ is not a vector space. Here, we outline an adaptation of the latter method.
Fix $\varepsilon\in(0,1)$, and define the frequency envelope of $f\in\mathcal{Z}^{2}_{\rho}$ by
$c_{j}[f]:=\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\|P_{k}\partial_{x}f\|_{H^{1}}+2^{-\varepsilon j}\big(\|f\|_{L^{\infty}}+\big\||f|^{2}-\rho^{2}\big\|_{L^{2}}^{\frac{1}{2}}\big),\qquad j\in\mathbb{Z}_{\geq 0}.$
@@
We collect some basic properties of $\mathbf{c}[f]:=\{c_{j}[f]\}_{j\geq 0}$ in the following lemma, which can be shown by a standard argument (see, e.g., [Ifrim-Tataru]); we will give a proof in Appendix B for completeness.
@@
For $f,g\in\mathcal{Z}^{2}_{\rho}$, the following holds.
@@ 1. $\mathrm{(i)}$
    $\|P_{j}\partial_{x}f\|_{H^{1}}\leq c_{j}[f]$ $(j\geq 0)$, $c_{k}[f]\leq 2^{\varepsilon|k-j|}c_{j}[f]$ $(j,k\geq 0)$, $\mathcal{E}^{2}_{\rho}(f)\sim\|\mathbf{c}[f]\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}^{2}$.
@@ 2. $\mathrm{(ii)}$
    $\big\|\mathbf{c}[f]-\mathbf{c}[g]\big\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}\lesssim\|\partial_{x}(f-g)\|_{H^{1}}+\|f-g\|_{L^{\infty}}+\big\||f|^{2}-|g|^{2}\big\|_{L^{2}}^{1/2}$.
@@ 3. $\mathrm{(iii)}$
    $\mathcal{E}^{3}_{\rho}(P_{\leq j}f)]^{1/2}\lesssim 2^{j}c_{j}[f]$ $(j\geq 0)$.
For any $\phi\in\mathcal{Z}^{2}_{\rho}$ with $\mathcal{E}^{2}_{\rho}(\phi)\leq M_{0}^{2}$ and $j\geq 1$, let $u$ and $u^{j}$ be the unique solutions of (gINLS) with initial condition, respectively, $u(0)=\phi$ and $u^{j}(0)=P_{\leq j}\phi$, which we have constructed in the class $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2}_{\rho})$.555 From (2.9), we see that $\mathcal{E}^{2}_{\rho}(P_{\leq j}\phi)\leq 3M_{0}^{2}$ for any $j\geq 0$. Therefore, by slightly modifying the choice of $T_{*}$ in Proposition 3.3 we can construct the solutions $u$, $u^{j}$ on a uniform interval $0,T_{*}]$. Note that $P_{\leq j}\phi\in\mathcal{Z}^{3}_{\rho}$ and hence $u^{j}\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{3}_{\rho})\subset C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$. Using the $\mathcal{Z}^{3}_{\rho}$ bound from (3.24), $\mathcal{Z}^{1}_{\rho}$ difference bound (3.26), and Lemma 3.4, it is not hard to derive the following estimate:
@@
This estimate is the key to obtain the continuity properties, showing that the initial frequency distribution is not significantly disturbed during the local existence time. For the reader’s convenience, we will also give a proof of (3.27) in Appendix B.
To see $u\in C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$, we first notice that the difference bound (3.26) shows the convergence $u^{j}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$, and hence $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho})$. It then suffices to prove $u\in C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$. We deduce from (3.27) that for any $m_{2}>m_{1}\geq 0$,
@@
$\lesssim\sum_{j_{1},j_{2}=m_{1}+1}^{m_{2}}2^{-\frac{1}{2}|j_{1}-j_{2}|}c_{j_{1}}[\phi]c_{j_{2}}[\phi]\lesssim\sum_{j=m_{1}+1}^{m_{2}}c_{j}[\phi]^{2}.$
Since $\mathbf{c}[\phi]\in\ell^{2}(\mathbb{Z}_{\geq 0})$, we see that the sequence $\{u^{j}\}_{j\geq 0}$ is Cauchy in $C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$, concluding that the limit $u$ belongs to $C([0,T_{*}];\mathcal{Z}^{2}_{\rho})$ and that
@@
To see the continuity of the flow map, let $\phi,\{\phi_{n}\}_{n\geq 0}$ be such that $\mathcal{E}^{2}_{\rho}(\phi),\mathcal{E}_{\rho}^{2}(\phi_{n})\leq M_{0}^{2}$ and $d^{2}_{\rho}(\phi_{n},\phi)\to 0$ as $n\to\infty$. Furthermore, let $u_{n},u_{n}^{j}\in C([0,T_{*}];\mathcal{Z}_{\rho}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}_{\rho}^{2})$ denote the solutions to (gINLS) with $u_{n}(0)=\phi_{n}$ and $u_{n}^{j}(0)=P_{\leq j}\phi_{n}$. In view of the $\mathcal{Z}^{1}_{\rho}$ difference estimate (3.26), it suffices to prove
@@
$\|\mathbf{c}[\phi_{n}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}\lesssim d^{2}_{\rho}(\phi_{n},\phi)+\big[d^{2}_{\rho}(\phi_{n},\phi)\big]^{\frac{1}{2}}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}.$
For the last term, we can use the $\mathcal{Z}^{3}_{\rho}$ bound from (3.24) for $u_{n}^{j}$ and $u^{j}$ as well as difference bound (3.26):
$\|\partial_{x}^{2}(u_{n}^{j}-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\leq\big(\|\partial_{x}^{3}u_{n}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{3}u^{j}\|_{L^{\infty}_{T_{*}}L^{2}}\big)\|d^{1}_{\rho}(u_{n}^{j},u^{j})\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{M_{0}}2^{j}\cdot d^{1}_{\rho}(P_{\leq j}\phi_{n},P_{\leq j}\phi)$
$\lesssim_{M_{0}}2^{j}\big(d^{1}_{\rho}(\phi_{n},\phi)+\|P_{\geq j+1}\phi_{n}\|_{L^{2}}+\|P_{\geq j+1}\phi\|_{L^{2}}\big)$
$\lesssim_{M_{0}}2^{j}d^{1}_{\rho}(\phi_{n},\phi)+2^{-j}.$
@@
$\limsup_{n\to\infty}\|\partial_{x}^{2}(u_{n}-u)\|_{L^{\infty}_{T_{*}}L^{2}}\lesssim_{M_{0}}\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}+2^{-\frac{1}{2}j}$

### v2

@@ 3. Local well-posedness
Let $\phi\in\mathcal{Z}^{2}$ satisfy $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$, and let $T_{*}>0$ be the constant given in Proposition 3.3. We divide the proof into several steps.
Step 1. We shall construct a local solution $u\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ to (gINLS) with initial data $\phi$ such that $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$ and
(3.24) $\big\|\mathcal{E}^{2}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim\mathcal{E}^{2}(\phi).$
We also show that if $\phi\in\mathcal{Z}^{k}$ for some integer $k\geq 3$, then $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{k})$ and it satisfies
(3.25) $\big\|\mathcal{E}^{k}(u)\big\|_{L^{\infty}_{T_{*}}}\lesssim_{k}\mathcal{E}^{k}(\phi).$
Let $\{u^{(\ell)}\}_{\ell\geq 0}$ be the sequence of solutions to the mollified problem (3.1) with $u^{(\ell)}(0)=P_{\leq\ell}\phi$ constructed in Proposition 3.1. We recall from Propositions 3.2 and 3.3 that $u^{(\ell)}\in C([0,T_{*}];\mathcal{Z}^{2})$ for all $\ell\geq 0$ and $\{u^{(\ell)}\}_{\ell\geq 0}$ is Cauchy in $C([0,T_{*}];\mathcal{Z}^{1})$. Since $C([0,T_{*}];\mathcal{Z}^{1})$ is a complete metric space, there exists $u\in C([0,T_{*}];\mathcal{Z}^{1})$ such that $u^{(\ell)}\to u$ strongly in $C([0,T_{*}];\mathcal{Z}^{1})$. Moreover, (3.5) shows that the sequence $\{\partial_{x}u^{(\ell)}\}_{\ell\geq 0}$ is bounded in $L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. By the Banach-Alaoglu theorem, there exists a subsequence converging weakly-* to some $\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$. Since $\partial_{x}u^{(\ell)}\to\partial_{x}u$ strongly in $C([0,T_{*}];L^{2}(\mathbb{R}))$, we have $\partial_{x}u=\widetilde{u}\in L^{\infty}(0,T_{*};H^{1}(\mathbb{R}))$, that is, $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{2})$. The bound (3.24) follows from (3.5) and weak-* lower semicontinuity of the $L^{\infty}_{T_{*}}H^{1}$ norm. The claim in the regular case $\phi\in\mathcal{Z}^{k}$ ($k\geq 3$) can be shown in the same way by using (3.6).
It remains to show that the limit $u$ is a solution to (gINLS) with $u(0)=\phi$ and belongs to $C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$. The initial condition is clearly satisfied by the strong convergence in $C([0,T_{*}];\mathcal{Z}^{1})$. To see that $u$ satisfies the integral equation
@@
$\lesssim 2^{-\ell}(1+\delta^{-1})\|\mathcal{E}^{2}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}^{\frac{3}{2}}+(1+\delta^{-1})\Big(1+\|\mathcal{E}^{1}(P_{\leq\ell}u^{(\ell)})\|_{L^{\infty}_{T_{*}}}+\|\mathcal{E}^{1}(u)\|_{L^{\infty}_{T_{*}}}\Big)\|d^{1}(P_{\leq\ell}u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{\delta,M_{0}}2^{-\ell}+\|d^{1}(u^{(\ell)},u)\|_{L^{\infty}_{T_{*}}}$
@@
where we have used Lemma 2.1, (2.10), (3.5), (3.24) and the strong convergence $u^{(\ell)}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1})$. In particular, we have the convergence of the Duhamel integrals
@@
This and the convergence of the linear part imply that $u$ satisfies (3.26). Since $U(t)\phi\in C(\mathbb{R};\mathcal{Z}^{1}_{\rho_{\phi}})$ by (2.17) and $\int_{0}^{t}U(t-s)\mathcal{N}(u(s))\,ds\in C([0,T_{*}];H^{1}(\mathbb{R}))$ for $u\in L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ by Lemma 2.1, we see from the integral equation (3.26) that $u\in C([0,T_{*}];\mathcal{Z}^{1}_{\rho_{\phi}})$.
@@
We conclude Step 1 by mentioning a bound of distance between two solutions of (gINLS). Let $\phi,\psi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi),\mathcal{E}^{2}(\psi)\leq M_{0}^{2}$, and let $u,v\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ be the corresponding solutions of (gINLS) constructed above. Then, taking $\ell,m\to\infty$ in (3.15), we see that
(3.27) $\big\|d^{1}(u,v)\big\|_{L^{\infty}_{T_{*}}}\lesssim(1+M_{0})d^{1}(\phi,\psi).$
@@
Step 2. We claim the following: Let $u,v\in C([0,T];\mathcal{Z}^{2})$ be two solutions to (gINLS) on an interval $[0,T]$. Assume that $\|\mathcal{E}^{2}(u)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ and $\|\mathcal{E}^{2}(v)\|_{L^{\infty}_{T}}\leq M_{1}^{2}$ for some $M_{1}>0$. Then, there exists $T_{1}=T_{1}(\delta,M_{1})>0$ such that for $0\leq t\leq\min\{T,T_{1}\}$ it holds that
$d^{1}(u(t),v(t))\lesssim(1+M_{1})d^{1}(u(0),v(0)).$
@@
In particular, the solution of the Cauchy problem associated with (gINLS) is unique in $C([0,T];\mathcal{Z}^{2})$.
@@
We remark that this claim must be discussed independently of the previous Step 1 (in particular the difference estimate (3.27)), since for uniqueness we have to consider general solutions which may differ from those constructed in Step 1. However, the claim can be proven along the same lines as Proposition 3.3 using the original equation (gINLS), (3.26) instead of the mollified one (3.1), (3.2). We omit the details.
@@
Step 3. We shall prove that the solution $u$ constructed in Step 1 satisfies $u\in C([0,T_{*}];\mathcal{Z}^{2})$ and the continuity of the flow map $u_{0}\mapsto u$ in $C([0,T_{*}];\mathcal{Z}^{2})$.
@@
We can prove $u\in C([0,T_{*}];\mathcal{Z}^{k})$ for $\phi\in\mathcal{Z}^{k}$ if $k\geq 3$ in a similar manner; we omit the proof. In particular, this concludes the proof of Theorem 1.1. To prove the claim, we may adapt an argument due to Bona–Smith [8] or the idea of frequency envelope due to Tao [42, 23, 3] to the non-vanishing solutions in $\mathcal{Z}^{2}$.444 The abstract theorem in [3] does not seem to apply directly to our problem, since $\mathcal{Z}^{2}$ is not a vector space. Here, we outline an adaptation of the latter idea.
Fix $\varepsilon\in(0,1)$, and define the frequency envelope of $f\in\mathcal{Z}^{2}$ by
$c_{j}[f]:=\sum_{k=0}^{\infty}2^{-\varepsilon|j-k|}\|P_{k}\partial_{x}f\|_{H^{1}}+2^{-\varepsilon j}\big(\|f\|_{L^{\infty}}+\big\||f|^{2}-\rho_{f}^{2}\big\|_{L^{2}}^{\frac{1}{2}}\big),\qquad j\in\mathbb{Z}_{\geq 0}.$
@@
We collect some basic properties of $\mathbf{c}[f]:=\{c_{j}[f]\}_{j\geq 0}$ in the following lemma, which can be shown by a standard argument (see, e.g., [23]); we give a proof in Appendix B for completeness.
@@
For any $f,g\in\mathcal{Z}^{2}$ and integers $j,k\geq 0$, the following hold.
@@ 1. $\mathrm{(i)}$
    $\|P_{j}\partial_{x}f\|_{H^{1}}\leq c_{j}[f]$, $c_{k}[f]\leq 2^{\varepsilon|k-j|}c_{j}[f]$, $\mathcal{E}^{2}(f)\sim\|\mathbf{c}[f]\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}^{2}$.
@@ 2. $\mathrm{(ii)}$
    $\big\|\mathbf{c}[f]-\mathbf{c}[g]\big\|_{\ell^{2}(\mathbb{Z}_{\geq 0})}\lesssim\|\partial_{x}(f-g)\|_{H^{1}}+\|f-g\|_{L^{\infty}}+\big\|(|f|^{2}-\rho_{f}^{2})-(|g|^{2}-\rho_{g}^{2})\big\|_{L^{2}}^{1/2}$.
@@ 3. $\mathrm{(iii)}$
    $\mathcal{E}^{3}(P_{\leq j}f)]^{1/2}\lesssim 2^{j}c_{j}[f]$.
For any $\phi\in\mathcal{Z}^{2}$ with $\mathcal{E}^{2}(\phi)\leq M_{0}^{2}$ and $j\geq 1$, let $u$ and $u^{j}$ be the unique solutions of (gINLS) with initial condition, respectively, $u(0)=\phi$ and $u^{j}(0)=P_{\leq j}\phi$, which we have constructed in the class $C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$.555 From (2.9), we see that $\mathcal{E}^{2}(P_{\leq j}\phi)\leq 3M_{0}^{2}$ for any $j\geq 0$. Therefore, by slightly modifying the choice of $T_{*}$ in Proposition 3.3 we can construct the solutions $u$, $u^{j}$ on a uniform interval $0,T_{*}]$. Note that $P_{\leq j}\phi\in\mathcal{Z}^{3}$ and hence $u^{j}\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{3})\subset C([0,T_{*}];\mathcal{Z}^{2})$. Using the $\mathcal{Z}^{3}$ bound from (3.25), the $\mathcal{Z}^{1}$ difference bound (3.27), and Lemma 3.4, we can derive the following estimate:
@@
This estimate is the key to prove the claim of Step 3 and shows that the initial frequency distribution is not significantly disturbed during the local existence time. For the reader’s convenience, we also give a proof of (3.28) in Appendix B.
To see $u\in C([0,T_{*}];\mathcal{Z}^{2})$, we first notice that the difference bound (3.27) shows the convergence $u^{j}\to u$ in $C([0,T_{*}];\mathcal{Z}^{1})$, and hence $u\in C([0,T_{*}];\mathcal{Z}^{1})$. It then suffices to prove $u\in C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$. We deduce from (3.28) that for any $m_{2}>m_{1}\geq 0$,
@@
$\lesssim\sum_{j_{1},j_{2}=m_{1}+1}^{m_{2}}2^{-\frac{1}{2}|j_{1}-j_{2}|}c_{j_{1}}[\phi]c_{j_{2}}[\phi]\lesssim\sum_{j=m_{1}+1}^{m_{2}}c_{j}[\phi]^{2}\leq\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>m_{1}})}^{2}.$
Since $\mathbf{c}[\phi]\in\ell^{2}(\mathbb{Z}_{\geq 0})$, we see that the sequence $\{u^{j}\}_{j\geq 0}$ is Cauchy in $C([0,T_{*}];\dot{H}^{2}(\mathbb{R}))$, concluding that the limit $u$ belongs to $C([0,T_{*}];\mathcal{Z}^{2})$ and that
@@
To see the continuity of the flow map, let $\phi,\{\phi_{n}\}_{n\geq 0}$ be such that $\mathcal{E}^{2}(\phi),\mathcal{E}^{2}(\phi_{n})\leq M_{0}^{2}$ and $d^{2}(\phi_{n},\phi)\to 0$ as $n\to\infty$. Furthermore, let $u_{n},u_{n}^{j}\in C([0,T_{*}];\mathcal{Z}^{1})\cap L^{\infty}(0,T_{*};\mathcal{Z}^{2})$ denote the solutions to (gINLS) with $u_{n}(0)=\phi_{n}$ and $u_{n}^{j}(0)=P_{\leq j}\phi_{n}$. In view of the $\mathcal{Z}^{1}$ difference estimate (3.27), it suffices to prove
@@
$\|\mathbf{c}[\phi_{n}]\|_{\ell^{2}(\mathbb{Z}_{>j})}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}\lesssim d^{2}(\phi_{n},\phi)+\big[d^{2}(\phi_{n},\phi)\big]^{\frac{1}{2}}+\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}.$
For the last term, we can use the $\mathcal{Z}^{3}$ bound from (3.25) and the difference bound (3.27) as well as (2.12):
(3.31) $\|\partial_{x}^{2}(u_{n}^{j}-u^{j})\|_{L^{\infty}_{T_{*}}L^{2}}^{2}$ $\leq\big(\|\partial_{x}^{3}u_{n}^{j}\|_{L^{\infty}_{T_{*}}L^{2}}+\|\partial_{x}^{3}u^{j}\|_{L^{\infty}_{T_{*}}L^{2}}\big)\|d^{1}(u_{n}^{j},u^{j})\|_{L^{\infty}_{T_{*}}}$
$\lesssim_{M_{0}}2^{j}\,d^{1}(P_{\leq j}\phi_{n},P_{\leq j}\phi)\lesssim_{M_{0}}2^{j}\,d^{1}(\phi_{n},\phi).$
@@
$\limsup_{n\to\infty}\|\partial_{x}^{2}(u_{n}-u)\|_{L^{\infty}_{T_{*}}L^{2}}\lesssim_{M_{0}}\|\mathbf{c}[\phi]\|_{\ell^{2}(\mathbb{Z}_{>j})}$